OpenStudy (anonymous):

@triciaal @perl Is the set of multiples of 6 closed under subtraction? Explain why or provide a counterexample if not. A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

OpenStudy (perl):

6 - 6 = 0 and 0 is a multiple of 6 , 6*0 = 0 6 - 12 = -6 and -6 is a multiple of 6 , 6*-1 = -6

OpenStudy (anonymous):

A

OpenStudy (perl):